Grade 12th passGeneral Physics

A body is thrown vertically upward with 45 m/sec distance travelled by the body in 5th second is(g=10m/s2)

Anirbna
7 Years agoGrade 12th pass

1 Answer

Mahima Kanawat
7 Years ago
Dear student
Here, Time to reach maximum height is u/g = 45/10 = 4.5 sec
Therefore in 5th second i.e., between 4 and 5 seconds, body reach to maximum height and return back.
As we know , distance travelled by body in last second of ascending is equal to first second of descending 
Distance travelled by body in first 0.5 seconds of descending is s = 1/2*10*(0.5)2 = 1.25 m
Therefore total distance is 1.25 m +1.25 m = 2.5 m
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT